OpenStudy (anonymous):

Solve using S=r(theta) Radians Central Angle Arc Length ? Pie over 3 radians 3/2 m

OpenStudy (anonymous):

I know im solving for 'r' but i just dont know how to do it.

OpenStudy (welshfella):

S = arc length, theta = pi/3 and r = ? 3/2 = r * pi/3 multiply both sides of the equation by 3/pi

OpenStudy (welshfella):

this will make the right hand side = r so the answer will be on left hand side

OpenStudy (anonymous):

Multiply by 3/pi or pi/3?

OpenStudy (welshfella):

by 3/pi

OpenStudy (welshfella):

because r * pi/3 * 3/pi = r

OpenStudy (anonymous):

Oooo ok thanks!

OpenStudy (welshfella):

then LHS will be 3/2 * 3/pi

OpenStudy (welshfella):